OpenStudy (anonymous):

derivative of ln[(x+1)/(x^3+1)] I know I use the quotient rule, but what do I do with the ln?

OpenStudy (zzr0ck3r):

so its the chain rule ln(x) is the outer function, so you will have \(\frac{1}{\frac{x+1}{x^3+1}}(\frac{x+1}{x^3+1})'\)

OpenStudy (zzr0ck3r):

@rysuh understand?

OpenStudy (anonymous):

Oh I see now! Thank you so much.

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

It's ugly.